OpenStudy (anonymous):

How do you find the slope-intercept form of ⅓y + x + ½y = -3

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }y+x+\frac{ 1 }{ 2 }y=-3\]You'll need to add the like terms of 1/3y and 1/2y:\[\frac{ 1 }{ 3 }+\frac{ 1 }{ 2 }=\frac{ 1 }{ 3 }*\frac{ 2 }{ 2 }+\frac{ 1 }{ 2 }*\frac{ 3 }{ 3 }\]\[\frac{ 2 }{ 6 }+\frac{ 3 }{ 6 }=\frac{ 5 }{6 }\] Therefore: \[x+\frac{ 5 }{ 6 }y=-3\] Now, get the y alone. Can you do the first step? :o)

OpenStudy (anonymous):

Is it \[y = \frac{ 5 }{ 6 }x + -3\] ?